# CONTROLLABLE LOCALIZED FACE ANONYMIZATION VIA DIFFUSION INPAINTING

### v1

@@ 1 Introduction
Traditional anonymization methods, such as blurring, pixelation, or masking, are simple but significantly degrade image quality and utility, and remain susceptible to reconstruction attacks [3]. Recent Generative Adversarial Network (GAN)-based [4] approaches generate more realistic results but often fail to preserve non-identifiable attributes [5, 6]. For example, FALCO [7] generates completely synthetic images that retain specific facial attributes but alter elements such as background, hairstyle, and accessories, as shown in Fig. 1(a).
@@ 2.2 Proposed Framework
Fig. 1: Overview of the proposed framework for face anonymization.
Noisy Latent Generation. We leverage the pre-trained encoder to encode the input image $x$ into a latent representation $z_{0}=\mathcal{E}(x)$. A forward diffusion process is then applied to $z_{0}$ over $T$ steps, producing the noisy latent $z_{T}$:
@@
Masked Latent Generation. To preserve background details and restrict edits to the facial region, we use a masked image as guidance. Specifically, we use a pre-trained face parsing model $f$ [14] to obtain a binary mask $M\in\{0,1\}$, where $M=1$ denotes the editable facial region, and $M=0$ denotes the preserved background. The masked image $x_{m}=x\odot(1-M)$ is then encoded into a latent code $c=\mathcal{E}(x_{m})$. which is concatenated with the noisy latent $z_{T}$ to guide the UNet, ensuring that only the masked region is regenerated.
@@
where $w_{t}=\lambda\sigma_{t}$, with $\lambda$ controlling the trade-off between target attribute enforcement and inpainting capability. $\sigma_{t}$ is adaptively decreased during the denoising process, allowing for stronger guidance in the early noisy steps and a weaker influence as the image becomes cleaner. Then we modify the reverse denoising process as:
@@
Input: real image $x$, synthesized target image $x_{tgt}$, $\lambda$ value, and number of denoising steps $T^{\prime}$

### v2

@@ 1 Introduction
Traditional anonymization methods, such as blurring, pixelation, or masking, are simple but significantly degrade image quality and utility, and remain susceptible to reconstruction attacks [Fantômas_2022_arXiv]. Recent Generative Adversarial Network (GAN)-based [4] approaches generate more realistic results but often fail to preserve non-identifiable attributes [8, 13]. For example, FALCO [1] generates completely synthetic images that retain specific facial attributes but alter elements such as background, hairstyle, and accessories, as shown in Fig. 1(a).
@@ 2.2 Proposed Framework
Fig. 1: The proposed framework for face anonymization.
Noisy Latent Generation. We leverage the pre-trained encoder $\mathcal{E}$ to encode the input image $x$ into a latent representation $z_{0}=\mathcal{E}(x)$. A forward diffusion process is then applied to $z_{0}$ over $T$ steps, producing the noisy latent $z_{T}$:
@@
Masked Latent Generation. To preserve background details and restrict edits to the facial region, we use a masked image as guidance. Specifically, we use a pre-trained face parsing model $f$ [12] to obtain a binary mask $M$ where $M_{i,j}\in\{0,1\}$, and $M_{i,j}=1$ denotes the pixel $(i,j)$ belongs to the editable facial region, and $M_{i,j}=0$ denotes the pixel belongs to the preserved background. The masked image $x_{m}=x\odot(1-M)$ is then encoded into a latent code $c=\mathcal{E}(x_{m})$. which is concatenated with the noisy latent $z_{T}$ to guide the UNet, ensuring that only the masked region is regenerated.
@@
where $w_{t}=\lambda\sigma_{t}$, with the guidance weight $\lambda$ controlling the trade-off between target attribute enforcement and inpainting capability. $\sigma_{t}$ is adaptively decreased during the denoising process, allowing for stronger guidance in the early noisy steps and a weaker influence as the image becomes cleaner. Then we modify the reverse denoising process as:
@@
Input: real image $x$, synthesized target image $x_{tgt}$, guidance weight $\lambda$, and number of denoising steps $T^{\prime}$
@@ 4 Conclusion
Acknowledgements. This work was supported by the Research Council of Finland (former Academy of Finland) Academy Professor project EmotionAI (grants 336116, 345122, 359894), HPC project FaceCanvas (grant number 364905), the University of Oulu & Research Council of Finland Profi 7 (grant 352788), EU HORIZON-MSCA-SE-2022 project ACMod (grant 101130271), Academy Research Fellow project (grant 355095), and Infotech Oulu.